An equally weighted portfolio consists of 12 assets which all have a standard deviation of 0.194. The average covariance between the assets is 0.155. Compute the standard deviation of this portfolio. Please enter your answer as a percentage to three decimal places (i.e. 12.345% rather than 0.12345 -- the percent sign is optional).
Answer #1

Variance of an equally weighted portfolio = (average variance of all assets in the portfolio/n) + [(n-1)/n]*average covariance between assets

= (0.194^2/12) + [(12-1)/12]*0.155 = 0.1452

Standard deviation = (0.1452)^0.5 = 38.108%